dear expert give an example of comutative law in real world.Best wishes
1
Expert's answer
2012-05-25T08:08:43-0400
Commutative Law
Is the Law that says you can swap numbers around and still get the same answer when you add. Or when you multiply.
& Examples:
& You can swap when you add: 3 + 6 = 6 + 3
& You can swap when you multiply: 2 × 4 = 4 × 2
In real world there are a lot of examples of commutative law. We meet it every day.
For example: in football. Coach can use different scheme of the game: 4-4-2; 5-3-1; 4-3-3 or other but only 10 field players can take part in the game.
Other example in the companies with different organize structure. Seniors can vary different positions and number of people in departments but in sum they have permanent quantity of personnel which is determined in company organize structure etc.